Allison:

4m^3/3m @blackghost356

Allison:

Rip

dude:

\(\frac{4m^3}{3m}\) They both have an m in common \(\frac{4m^{3\color{red}{-1}}}{3m^{1\color{red}{-1}}}\rightarrow \frac{4m^2}{3\times 1}\rightarrow \boxed{\frac{4m^2}{3}}\)

Allison:

Idonyunderstandanyofthisbutyaywhocaresitsstupidalgebrahehethankyoubebeh

Allison:

dOnY

Allison:

@dude

Allison:

Enuff